10. (**) [2 points A firm solves the cost minimization problem and finds the follow- ing: The isocost is tangent to the isoquant (of q 100) such that K-100 and l. : 였). Ille íirm know: ; íhai, r X) and w :X. What is the cosi of producing 100 units of output? Show how you got the answer
Answer #1

The cost of producing 100 units of output is given as:

C = w*L + r*K

= 7*20 + 20*100 = 2140.

Thanks!